OpenStudy (aykayyy):

solve for x here 6+2x+(3x^2)/(1-x)<3/(1-x)

OpenStudy (jack1):

\[\large \frac{(6+2x+(3x^2)}{(1-x)}<\frac{3}{(1-x)}\] or \[\large 6+2x+\frac{(3x^2)}{(1-x)}<\frac{3}{(1-x)}\] ...which?

OpenStudy (aykayyy):

the bottom one

OpenStudy (anonymous):

Your answer is \[-1<x<3\]

OpenStudy (aykayyy):

the answer is x>3 but i dont understand why... when i do it i get x<3

OpenStudy (anonymous):

Sorry, I think I made a mistake. It should be \[1 < x <3\] I don't see why it should be x>3.

OpenStudy (aykayyy):

i dont either... which is why im confused lol

OpenStudy (anonymous):

Hmmm...

OpenStudy (jack1):

6+2x+((3x^2))/(1-x) < 3/(1-x) 6(1-x)+2x(1-x)+(3x^2) < 3 6-6x + 2x-2x^2 + 3x^2 < 3 6 - 6x +2x -2x^2 + 3x^2 < 3 6 - 4x + x^2 < 3 3 - 4x + x^2 < 0 solve the quadratic 3 - 4x + x^2 = 0 (x-3)(x-1) = 0 therefore solutions are: x = 3 or x = 1 sub back in and test solutions: if x = 1: 6+2x+((3x^2))/(1-x) < 3/(1-x) 6+2+((3))/(1-1) < 3/(1-1) 11/(0) < 3/(0) cant divide by zero, so 1 isn't a solution now test other solution: if x = 3 6+2x+((3x^2))/(1-x) < 3/(1-x) 6+6+((27))/(1-3) < 3/(1-3) 39/-2 < 3/-2 -18.5 < -1.5 true, so 3 is the smallest solution x>3

OpenStudy (jack1):

sorry, i type really slowly, cheers for waiting

OpenStudy (jack1):

i guess if u want to "prove" the maths, u can test any number smaller than 3 and any number larger than 3... it should hold up

OpenStudy (anonymous):

Hurray!

OpenStudy (aykayyy):

thank you @jack1